### v1

# NICER Magnetar Burst Catalog

Abstract
In this paper, we present a comprehensive catalog of short bursts from magnetars based on eight years of NICER observations. A total of 1130 bursts were identified, making this the largest magnetar burst catalog to date. The sample is dominated by SGR 1935+2154, which contributes 76% of all detected bursts. We analyzed burst durations, spectral properties, and their correlations across multiple sources. Bursts from SGR 1935+2154 exhibit significantly longer durations, with a mean of 317 ms, compared to a mean of 23 ms for bursts from other magnetars. Two s-scale bursts were detected for the first time, originating from 1E 1048.15937 and CXOU J010043.1721134. Spectral analysis in the 0.5–8 keV range using both blackbody and power-law models shows that bursts with higher fluences have harder spectra. In contrast, correlations between burst duration and spectral parameters are weak or absent. This catalog provides a valuable dataset for studying magnetar short bursts, enabling future modeling efforts and improving our understanding of the diversity and physical mechanisms of magnetar bursts.
1 Introduction
Magnetars are a class of neutron stars characterized by their extremely strong magnetic fields, often exceeding – G (1992ApJ...392L...9D). In the X-ray band, unlike rotation-powered pulsars, which typically show steady emission, magnetars are known for their highly variable and transient behavior (e.g., 2018MNRAS.474..961C). The first discovered magnetar burst was the giant flare from SGR 052666 in 1979, followed by a short burst (1979Natur.282..587M). Shortly afterwards, a series of short bursts was detected from SGR 1900+14 (1979SvAL....5..343M). Because these repeating events are dominated by emission in the soft -ray band, the sources were named soft gamma repeaters (SGRs). During a magnetar outburst, which can last from several months to several years, short bursts often appear in clusters known as burst forests or burst storms, with rates that can reach hundreds of bursts per hour (e.g., 2020ApJ...904L..21Y; 2025ApJ...989...63H). In contrast, during the quiescent state, bursts may appear individually (e.g., 2002Natur.419..142G). According to the magnetar model, these bursts are generally thought to be triggered by cracks in the neutron star crust, induced by extreme magnetic stress (1992ApJ...392L...9D; 1995MNRAS.275..255T).
With the discovery of more SGRs, several short-burst catalogs of individual sources have been reported, including SGR 180620, SGR 1900+14, 1E 1547.05408, SGR 0501+4516, etc. (2001ApJ...558..228G; 2012ApJ...749..122V; 2013ApJ...778..105L). Most of these studies focused on bursts in the hard X-ray band (10 keV), while some extended to the soft X-ray range (10 keV). The burst durations are typically log-normally distributed over more than two orders of magnitude, peaking around 200 ms. Their X-ray spectra can generally be described equally well by models such as a double blackbody, optically thin thermal bremsstrahlung, or a Comptonized model. A clear correlation between spectral hardness and brightness has been observed in these catalogs, where bursts with higher fluence tend to exhibit harder spectra, regardless of the adopted spectral model (2001ApJ...558..228G; 2012ApJ...749..122V; 2013ApJ...778..105L).
A more comprehensive burst catalog was later compiled using data from Fermi/Gamma-ray Burst Monitor (GBM), which included bursts from more than eight magnetar sources, though the majority originated from 1E 1547.05408 (2015ApJS..218...11C). The timing and spectral properties reported in this catalog are consistent with those obtained from earlier observations. In the 2020s, several magnetars underwent new outbursts that produced numerous short bursts. Detailed analyses have been conducted for bursts from Swift J1555.25402, SGR 18300645, and SGR 1935+2154 (2021ApJ...920L...4E; 2022ApJ...924..136Y; 2025ApJ...989...63H). While their general properties are similar to those of other magnetar bursts, the events from SGR 1935+2154 stand out with substantially longer durations, peaking around 1 s, which is significantly longer than other magnetars (2020ApJ...904L..21Y).
The longer-duration short bursts were observed during the 2020 outburst of SGR 1935+2154. During this outburst, a fast radio burst (FRB) 20200428, was detected and confirmed to be associated with a short burst from SGR 1935+2154 (2020Natur.587...54C; 2020Natur.587...59B; 2020ApJ...898L..29M). This discovery established SGR 1935+2154 as the only Galactic magnetar known to emit FRBs. The FRB 20200428 associated X-ray short burst has a similar duration to other short bursts from this source (2021NatAs...5..378L), however, its spectrum is significantly softer (2021NatAs...5..408Y). Beyond the distinctions seen within SGR 1935+2154 itself, 2013ApJ...778..105L reported systematic differences in duration distributions between magnetars. Together, these findings highlight the diversity of magnetar burst behavior, suggesting that physical conditions and emission mechanisms vary among sources and outbursts. A systematic comparison of bursts from multiple magnetars is therefore essential to constrain the physical origin and diversity of magnetar short bursts.
In this paper, we present a systematic study of magnetar short bursts using eight years of Neutron Star Interior Composition Explorer (NICER) observations. We compile a catalog of bursts from multiple magnetars, including the highly active SGR 1935+2154 as well as magnetar-like pulsars (2008Sci...319.1802G; 2016ApJ...829L..21A; 2016ApJ...829L..25G), and perform analysis of their timing and spectral properties. We also investigate correlations among these properties and explore variations between different sources. This comprehensive analysis provides a uniform catalog of magnetar short bursts observed by NICER, offering a resource for future modeling and theoretical studies.
| Magnetars | Number of | Exposure |
| observations | (ks) | |
| CXOU J010043.1721134 | 241 | 725.0 |
| 4U 0142+61 | 166 | 330.0 |
| SGR 0418+5729 | 16 | 21.3 |
| SGR 0501+4516 | 198 | 498.9 |
| 1E 1048.15937 | 319 | 383.2 |
| 1E 1547.05408 | 60 | 87.9 |
| Swift J1555.25402 | 68 | 118.1 |
| PSR J16224950 | 22 | 46.0 |
| CXOU J164710.2455216 | 111 | 166.9 |
| 1RXS J170849.0400910 | 113 | 84.9 |
| CXOU J171405.7381031 | 1 | 3.7 |
| SGR J17452900 | 55 | 91.9 |
| SGR 180620 | 15 | 37.2 |
| XTE J1810197 | 168 | 514.2 |
| Swift J1818.01607 | 74 | 181.2 |
| Swift J1822.31606 | 12 | 14.7 |
| SGR 18300645 | 78 | 245.1 |
| 1E 1841045 | 170 | 386.6 |
| SGR 1900+14 | 1 | 0.4 |
| SGR 1935+2154 | 415 | 1094. |
| 1E 2259+586 | 149 | 179.7 |
| PSR J11196127∗ | 111 | 157.0 |
| PSR J18460258∗ | 202 | 648.9 |
| Total | 2765 | 6016.9 |
| ∗ Magnetar-like pulsars | ||
2 Data and Analysis
2.1 Data Selection and Reduction
NICER is a soft X-ray (0.2–12 keV) telescope mounted on the International Space Station (2016SPIE.9905E..1HG). Its X-ray Timing Instrument (XTI) consists of 56 Focal Plane Modules (FPMs), though only 52 modules were in operation. While it does not offer imaging capability, NICER provides exceptionally high time resolution ( 100 ns) and a large effective area, making it ideal for studying transient events. For this work, we selected all NICER observations conducted from its launch, 2017 June, through 2025 June 30 that included a magnetar or a magnetar-like pulsar within the field of view. Among the 26 known magnetars (2014ApJS..212....6O)111McGill Online Magnetar Catalog http://www.physics.mcgill.ca/~pulsar/magnetar/main.html, NICER has observed 21 of them, along with two additional magnetar-like pulsars. We collected 6016.9 ks of exposure time from 2765 NICER observations across 23 sources. The total exposure time and number of observations used for each source are summarized in Table 1.
We used heasoft version 6.34 (2014ascl.soft08004N) to reduce the NICER data. The standard calibration and screening were performed by the task nicerl2 with default options. The task barycorr was then applied with the JPLEPH.430 ephemeris to calibrate the time to Barycentric Dynamical Time (TDB). The resulting event lists were separated into the standard (0.5–8 keV) and high-energy (10–12 keV) bands for magnetar burst searches. On 2023 May 22, NICER sustained damage to its optical blocking film, leading to optical light leakage that degraded data quality. Observations taken after this date were processed separately. Nighttime data were reduced using the standard procedure described above. For daytime data, the nicerl2 task was executed with a screening option of threshfilter=DAY to account for increased optical contamination.
| Sky region | ObsID | Exposure (ks) | Counts |
| RXTE1 | 6012010106 | 3.1 | 1916 |
| RXTE2 | 1012020169 | 18.6 | 6924 |
| RXTE3 | 1012030123 | 5.3 | 2812 |
| RXTE4 | 5012040217 | 8.3 | 4336 |
| RXTE5 | 1012050131 | 5.8 | 4234 |
| RXTE6 | 4012060255 | 13.3 | 7400 |
| RXTE8 | 1012080190 | 12.2 | 5574 |
2.2 Burst Searching
We applied the Bayesian block algorithm to both the standard and high-energy event lists to identify significant flux variations (1998ApJ...504..405S). Blocks occurring during high-particle-background intervals, as defined by the high-energy band, were excluded. In the standard band, blocks with durations shorter than 1 s were selected as burst candidates. We further examined the FPM distribution of each candidate. Candidates with events concentrated in only one or two FPMs were identified as particle-induced events and excluded from the burst sample.
For each burst candidate, we estimated the statistical significance by computing the Poisson probability of detecting the observed number of counts within the burst block. A nearby interval of 5 s was selected to represent the non-burst count rate. After accounting for the number of trials, we retained only those blocks with a Poisson probability less than for further analysis. The number of trials was defined as the total exposure time of each observation divided by 0.2 s, which is a typical magnetar burst duration (2015ApJS..218...11C).
NICER has been monitoring seven X-ray blank regions defined initially by the Rossi X-ray Timing Explorer (RXTE; 2006ApJS..163..401J). We selected one observation from each region and excluded intervals with high particle background. The details of the blank sky observations used are summarized in Table 2. We binned the cleaned event lists into 0.2 s and 0.02 s bins, matching the burst durations reported in previous work (2015ApJS..218...11C) as well as those found in this study. We then recorded the total counts in each bin to characterize the background X-ray distribution. The cumulative distribution function reaches 100% at 4 counts per 0.2 s bin and 3 counts per 0.02 s bin (Figure 1). To ensure robust separation between genuine bursts and background fluctuations, we required that a burst candidate contain at least 5 counts within a Bayesian block. Based on the measured background rate and the NICER effective area, the expected detection limits for a transient event are approximately 2 and 2.5 erg s-1 cm-2 for transient duration 0.02 s and 0.2 s respectively.
| Magnetars | Number of burst |
| CXOU J010043.1721134 | 4 |
| 4U 0142+61 | 4 |
| SGR 0418+5729 | 0 |
| SGR 0501+4516 | 0 |
| 1E 1048.15937 | 1 |
| 1E 1547.05408 | 0 |
| Swift J1555.25402 | 74 |
| PSR J16224950 | 0 |
| CXOU J164710.2455216 | 0 |
| 1RXS J170849.0400910 | 1 |
| CXOU J171405.7381031 | 0 |
| SGR J17452900 | 0 |
| SGR 180620 | 4 |
| XTE J1810197 | 1 |
| Swift J1818.01607 | 27 |
| Swift J1822.31606 | 1 |
| SGR 18300645 | 128 |
| 1E 1841045 | 15 |
| SGR 1900+14 | 0 |
| SGR 1935+2154 | 865 |
| 1E 2259+586 | 1 |
| PSR J11196127∗ | 0 |
| PSR J18460258∗ | 4 |
| Total | 1130 |
| ∗ Magnetar-like pulsars | |
2.3 Analysis
We first plotted the cumulative counts of each selected burst, after subtracting the non-burst count rate, and fitted the resulting curve with a step function (2001ApJ...558..228G). The step height corresponds to the total number of counts in the burst. This total count was then used to compute the 5% and 95% levels of the cumulative distribution. The time interval between when the cumulative count reaches 5% and 95% is defined as the duration (1993ApJ...413L.101K). For bursts with fewer than 20 total counts, the 5% fractional level corresponds to less than one count, making not accurately defined. In these cases, is instead defined as the duration of the corresponding Bayesian block.
The Bayesian blocks interval of each burst defines its good time interval (GTI), and the counts within this GTI were used for the spectral analysis. We first calculated the hardness ratio (HR), defined as
where soft denotes photons in the 0.5–4 keV range and hard denotes photons in the 4–8 keV range. Some bursts show no photons below 2 keV because of strong absorption along the line of sight, so we adopted 4 keV as the boundary between the soft and hard photons, rather than 2 keV.
We also extracted 0.5–8 keV burst spectra from the burst GTI with the task nicerl3-spect and performed spectral fitting using xspec v12.14 (2014ascl.soft08004N). The non-burst interval used for estimating the burst significance was also adopted as the background spectrum for subtraction during fitting. The spectra were binned according to the total number of burst counts: for bursts with fewer than 20 counts, at least 1 count per bin; for 20–49 counts, each bin contained at least 5 counts; for 50–249 counts, at least 10 counts per bin; and for more than 250 counts, at least 25 counts per bin. Each burst spectrum was fitted with both a blackbody and a power-law model. In all fits, the hydrogen column density () was fixed to values reported in the literature for each of the magnetars (2023MNRAS.526.1287C; 2021ApJ...920L...4E; 2021MNRAS.504.5244B; 2020ApJ...902....1H; 2022ApJ...924..136Y; 2023ApJ...952..120H). For bursts with fewer than 20 total counts, the blackbody temperature () or the power-law photon index () was also fixed at 1.2 keV and 0.8, respectively. The cstat in the xspec statistic was used for bursts with fewer than 250 counts, while the chi statistic was adopted for bursts with 250 counts or more. Once the best-fit parameters were obtained, we calculated the absorbed flux for both models in the 0.5–10 keV range. All uncertainties reported in this paper correspond to 1 confidence levels.
3 Results and Discussion
3.1 Overall samples
We detected a total of 1130 bursts in our dataset, including four from the magnetar-like pulsar PSR J18460258. The majority of events (76.5%) originated from SGR 1935+2154, which has been the most active magnetar in recent years. NICER captured its major outbursts in April 2020, September 2021, and October 2022 (2020ApJ...904L..21Y; 2021ATel14916....1G; 2025ApJ...989...63H), during which some Galactic FRBs were also observed (2020Natur.587...54C; 2020Natur.587...59B; 2021NatAs...5..414K; 2022ATel15681....1D; 2023arXiv231016932G).
Several magentar outbursts have also been captured by NICER since its launch in 2017 June, including the 2020 outburst of Swift J1818.01607 (2020ApJ...902....1H), the 2020 outburst of SGR 18300645 (2022ApJ...924..136Y), the 2021 outburst of Swift J1555.25402 (2021ApJ...920L...4E), and the 2024 outburst of 1E 1841045 (2025ApJ...989...89Y). These sources accounted for 21.6% of the bursts in our sample. NICER also observed additional outbursts from other magnetars, such as XTE J1810197 (2019ApJ...877L..30G), and PSR J18460258 (2023ApJ...952..120H). However, in these cases, the outburst activity was relatively weak or the observations did not coincide with the outburst onset, resulting in fewer detected bursts.
The number of bursts detected from each source is summarized in Table 3, while details of individual bursts are provided in Appendix A. Several of these bursts have also been reported in earlier studies focused on individual sources, and here we place them in the broader context of the NICER magnetar sample.
3.2 duration
The distribution of durations for all bursts is shown in Figure 2. The overall sample has a mean of 350 ms and a standard deviation of 470 ms. Two s-scale bursts were identified in our sample: one from 1E 1048.15937 with a duration of 0.6 s, and another from CXOU J010043.1721134 with a duration of 5 s. These represent the first detections of s-scale bursts in X-rays from magnetars. Their light curves are shown in Figure 3. Since no sub-ms bursts were found, it remains unclear whether these events are simply the extreme tail of the main burst distribution or instead represent a distinct class of magnetar short bursts. Future X-ray missions with larger effective area will be crucial to resolve this question.
Bursts from SGR 1935+2154 exhibit values that are significantly longer than those of other magnetars. We fitted log-normal distributions to both groups and show the normalized distributions in Figure 2. For SGR 1935+2154, the distribution peaks at 316 ms, while it peaks at 23 ms for the remaining magnetars. Thus, bursts from SGR 1935+2154 are on average about an order of magnitude longer. A Kolmogorov–Smirnov (KS) test applied to the two distributions yields a probability of , confirming that the two samples are distinct.
| Magnetars | Instrument | peak | 16–84% quantile |
| (ms) | (ms) | ||
| SGR 1935+2154 | NICER/XTI | 317 | 95–1026 |
| Other magnetars | NICER/XTI | 23 | 10–52 |
| SGR 1935+21541 | NICER/XTI | 840 | 430–1630 |
| 1E 1547.054082 | Swift/XRT | 207 | 62–685 |
| 1E 1547.054083 | Fermi/GBM | 156 | 62–389 |
| References. (1) 2020ApJ...904L..21Y (2) 2013ApJ...778..105L | |||
| (3) 2015ApJS..218...11C | |||
2020ApJ...904L..21Y studied bursts from SGR 1935+2154 during its 2020 outburst using NICER data. These events are also included in our sample, accounting for roughly 40% of the SGR 1935+2154 bursts analyzed here. However, the durations we obtain are systematically shorter than those reported in 2020ApJ...904L..21Y. A key reason is the different treatment of multi-peak bursts: we apply a 0.2 s criterion to decide whether peaks belong to separate bursts, whereas their study used 0.5 s. Consequently, they classified more events as multi-peak bursts, yielding shorter durations if treated as individual bursts. A second factor is the difference in burst populations across epochs. NICER captured a burst storm from SGR 1935+2154 in the 2020 outburst, during which the persistent emission was higher than usual. Dim or very short bursts may have been buried in this elevated background and thus missed. In contrast, the bursts detected in the 2021 and 2022 outbursts occurred during lower persistent-emission states, making weaker and shorter events easier to identify. As a result, the 2021 and 2022 bursts exhibit shorter duration distributions overall, pulling the combined distribution toward smaller values.
Regardless of methodology, the burst durations of SGR 1935+2154 differ significantly from those of other magnetars, suggesting possible intrinsic differences between this source and the broader magnetar population. SGR 1935+2154 is currently the only Galactic magnetar known to produce FRBs (2020Natur.587...54C; 2020Natur.587...59B), and numerous studies have explored what physical mechanisms might make it distinct. 2025ApJ...989...63H reported that during the 2022 outburst of SGR 1935+2154, the burst properties evolved across different epochs, with a notable change after an intermediate flare. The X-ray bursts that occurred after the flare, including those around the time of FRB 20221014A (2022ATel15681....1D), exhibit properties that differ significantly from the pre-flare burst population. 2021NatAs...5..408Y also found that the X-ray counterpart of FRB 20200428 exhibited a softer spectrum compared to typical bursts, potentially indicating an origin in a different region of the magnetosphere.
These findings suggest that the differences in durations between SGR 1935+2154 and other magnetars may reflect variations in the burst mechanism across different outbursts and even between epochs within a single outburst. Our SGR 1935+2154 sample combines bursts from the 2020, 2021, and 2022 outbursts, and further disentangling these epoch-dependent effects will be necessary to understand the underlying physics. A detailed comparison of burst properties across different epochs of SGR 1935+2154 will be presented in a forthcoming paper (Chu et al., in preparation). Together, these results emphasize that SGR 1935+2154 is an exceptional case among magnetars, potentially connected to its unique ability to produce FRBs.
In addition to bursts from other magnetars, 2013ApJ...778..105L reported 275 bursts from 1E 1547.05048 detected with Swift/X-ray Telescope (XRT), with a distribution peaking at 207 ms and 262 bursts from SGR 0501+4516 detected with XMM-Newton, whose durations peak at about 100 ms. 2015ApJS..218...11C presented a catalog of 440 bursts from at least eight magnetars observed by Fermi/GBM. Among these, 386 bursts originate from 1E 1547.05048 and have a peak at 156 ms, while the remaining bursts peak around 80 ms. Overall, these duration distributions are longer than those of the regular magnetar bursts in our NICER sample. Because NICER offers superior time resolution and larger effective area than those aforementioned missions, it is more sensitive to faint or shorter duration bursts. As a result, our catalog contains many more short events, leading to systematically shorter distributions compared to previous studies. A summary of the log-normal fitting results of the distributions, including the peak values and 16–84% quantile ranges for different groups and instruments, is presented in Table 4.
3.3 Hardness ratio
The distribution of the HR is shown in Figure 4. The overall distribution has a mean of 0.29 and a standard deviation of 0.31. Most bursts exhibit relatively soft spectra, while those from Swift J1555.25402 and Swift J1818.01607 are noticeably harder, with HR mostly greater than zero. This trend is consistent with the higher column densities of these sources ( and atoms cm-2, respectively), which preferentially absorb more soft photons compared to other magnetars.
For SGR 1935+2154, the HR distribution is well described by a normal distribution with a mean of 0.30 and a standard deviation of 0.21. This distribution is similar to those of other low- magnetars, in contrast to the marked differences observed in their burst-duration distributions.
3.4 Spectral model
The spectra of magnetar bursts in the soft X-ray band can be equally well described by either a single power-law or a single blackbody model (e.g., 2017ApJS..232...17K). Therefore, each burst spectrum was fitted individually with both models to account for both possibilities.
The results of the blackbody fits are shown in Figure 5, which presents the distributions of blackbody temperature () and flux (). Fitting the distribution of SGR 1935+2154 group with a log-normal function yields a peak at 1.56 keV and a 16–84% quantile range of 1.17–2.08 keV. This characteristic temperature is slightly lower than that obtained from Swift/XRT bursts of 1E 1547.05408 (2.5 keV; 2013ApJ...778..105L) and considerably lower than that of Fermi/GBM bursts (4.5 keV; 2015ApJS..218...11C). The difference likely reflects NICER’s higher sensitivity to fainter bursts, which tend to exhibit softer spectra (2001ApJ...558..228G; 2013ApJ...778..105L), corresponding to lower blackbody temperatures (see also Section 3.5). As a result, the inclusion of additional dim bursts shifts the overall distribution toward lower values. The higher temperatures inferred from Fermi/GBM may also arise because those fits correspond to the lower-temperature component of a two-blackbody model in the 8–200 keV range, where the cooler component is poorly constrained and thus biased upward. Alternatively, the broadband spectra may be better represented by three blackbody components rather than two. In any case, the burst temperatures are significantly higher than the typical magnetar temperature in quiescence of 0.45 keV (2023MNRAS.526.1287C), and even during the onset of outburst of 0.8 keV (e.g. 2019ApJ...874L..25G). The fluxes are generally distributed between and erg s-1 cm-2, with the two s-scale bursts showing much higher fluxes of erg s-1 cm-2.
The results of the power-law spectral fits are shown in Figure 6, which display the distributions of photon index () and corresponding flux (). A normal distribution fit to SGR 1935+2154 group yields a mean of 0.45 and a standard deviation of 0.44. Overall, the burst spectra are systematically harder than those of the magnetars persistent emission in quiescence. The flux distribution obtained from the power-law model is broadly consistent with that derived from the blackbody model, indicating that the inferred burst fluxes are not strongly dependent on the chosen spectral model.
The fluence ( and ) distributions, calculated from the fluxes and durations, for both the blackbody and power-law models are shown in Figure 7 as log()–log() plots. For fluences above erg cm-2, the distributions are well described by power-laws, yielding indices of 0.53 0.04 for the blackbody model and 0.52 0.03 for the power-law model. Because the full sample is dominated by SGR 1935+2154, we also fit the distributions after excluding its bursts. This yields significantly steeper slopes of for the blackbody and for the power-law model. These results demonstrate that SGR 1935+2154 exhibits a distinct fluence distribution compared to other magnetars. Its longer burst durations lead to a larger number of high-fluence events, which strongly influences the overall distribution.
From the turnover in these distributions, we estimate that NICER achieves nearly 100% detection efficiency above erg cm-2, with a detection limit of approximately erg cm-2. For the duration distribution of our burst sample, this detection limit is slightly higher than the expected value of erg cm-2 estimated in Section 2.2. The reason is that most of the bursts identified in this work occurred during a magnetar X-ray outburst, where the persistent emission is higher. As a result, the background level is higher, and some low-fluence bursts are buried in the background.
3.5 Correlation
The left panels of Figure 8 show the distributions of and versus their corresponding fluxes. For low-count bursts, where or were fixed during fitting, these points are not included in the figure. Overall, bursts with higher fluxes tend to exhibit harder spectra, characterized by higher temperatures or smaller photon indices, consistent with previous studies (see, e.g., 2001ApJ...558..228G; 2013ApJ...778..105L). To quantify this relation, we calculated Pearson correlation coefficients (Pearson r) for bursts from SGR 1935+2154 and from other magnetars separately. The clear outlier, the 5 s burst from CXOU J010043.1721134, was excluded from this analysis.
For SGR 1935+2154, the coefficients are 0.754 for versus and 0.702 for versus . For the other magnetar bursts, the corresponding values are 0.807 and 0.556, respectively. Strong correlations are found in all cases, except for the weaker – relation in the other magnetar sample. We fitted power-law functions to these highly correlated relations, obtaining indices of 0.18 0.01 and 0.22 0.02 for SGR 1935+2154 and the other magnetars, respectively, for the - relation. For the - relation, the indices are 0.51 0.03 for SGR 1935+2154 and 0.42 0.08 for the remaining magnetars. In addition to the s burst, two other events exhibit notably soft spectra ( keV or ): one from SGR 1935+2154 and one from 4U 0142+61.
The right panels of Figure 8 show the distributions of and versus and , respectively. Similar to the flux relations, bursts with higher fluences generally exhibit harder spectra. For SGR 1935+2154, the Pearson correlation coefficients are 0.588 for versus and for versus , while for the other magnetar bursts, they are 0.574 and , respectively. The correlations weaken when moving from flux to fluence, as the scatter increases at higher fluences, suggesting the presence of possible subgroups. However, after incorporating the duration parameter, the 5 s burst aligns with the overall trend, remaining consistent with the main burst population.
We also examined the fluence relations using power-law fits. For the - relation, the derived indices are 0.13 0.01 for SGR 1935+2154 and 0.19 0.02 for the other magnetars. In the - relation, the indices are 0.32 0.03 for SGR 1935+2154 and 0.28 0.07 for the remaining sources. Because these fits are driven primarily by the high-fluence end of the distribution, the results appear weaker and less well defined than those seen in the flux relations.
To further investigate possible subgroups within our sample, we performed a bootstrap analysis of the spectral correlations. For each selected burst sample, we generated a resampled dataset by randomly selecting bursts from the original sample until the resampled dataset matched the original in size. Each burst could be selected multiple times. We then recalculated the Pearson correlation coefficient for each resampled dataset. This process was repeated one million times, and the resulting distributions of correlation coefficients are shown in Figure 9.
The mean correlation coefficients derived from the bootstrap analysis are consistent with those obtained from the original samples across all parameter combinations. However, the coefficient distributions for the other magnetar bursts are noticeably broader than those for the SGR 1935+2154 sample, particularly for the power-law model. This wider spread likely reflects the inclusion of bursts from multiple magnetars, whose spectral parameters may be influenced by instrumental limitations and varying absorption along different lines of sight. The bootstrap results show no additional outlying behavior, and no secondary clusters or anomalous distributions appear across the one million trials. This suggests that all bursts follow a similar trend, except for the 5 s burst in the flux relation.
The correlations between burst duration and spectral parameters are shown in Figure 10, along with their bootstrap distributions. As with the spectral–spectral relations, bursts were divided into two groups: SGR 1935+2154 and all other magnetars. The two s-scale bursts were excluded from the analysis. For SGR 1935+2154, all correlations between and the spectral parameters are statistically insignificant, with Pearson coefficients consistent with zero. This is consistent with earlier studies of magnetar bursts, which also found no clear correlation between timing and spectral properties (2012ApJ...749..122V).
In contrast, for the other magnetars, both and show an apparent anti-correlation with . This trend is not intrinsic but instead results from instrumental limitations. Bursts with lower fluence are more likely to fall below the detection threshold, truncating the intrinsic burst population. The observed anti-correlation is therefore a selection effect. Similarly, the weak correlations and anti-correlations between and spectral hardness, represented by HR, , and , can be explained by the same detection bias. Shorter bursts tend to have lower fluxes, and as shown in Figure 9, fainter bursts generally display softer spectra. This leads to the weak correlations between duration and spectral parameters observed here.
4 Summary
In this study, we constructed a comprehensive catalog of magnetar short bursts spanning eight years of NICER observations. We identified a total of 1130 bursts, making this the largest burst catalog to date. Most of the bursts originated from SGR 1935+2154, which contributed 865 events. We present the distributions of timing and spectral parameters for all detected bursts and investigate their correlations. The bursts from SGR 1935+2154 generally show longer durations than those from other magnetars. We also report the first detection of s-scale bursts in magnetars, one from 1E 1048.15937 and another from CXOU J010043.1721134.
Bursts with higher fluxes tend to exhibit harder spectra, indicating higher blackbody temperatures in the blackbody model or lower photon indices in the power-law model. The bootstrap resampling analysis confirms that these spectral correlations are statistically robust, while correlations between burst duration and spectral parameters remain weak or absent. The weaker correlations found among other magnetars likely reflect differences in source properties and instrumental effects.
This catalog provides a uniform, statistically rich sample that will serve as a foundation for future investigations of magnetar burst energetics, emission mechanisms, and magnetic field dynamics.
Appendix A The Catalog
This appendix presents all bursts identified in this study. Magnetars for which at least one burst was detected include CXOU J010043.1721134 (Table 5), 4U 0142+61 (Table 6), 1E 1048.15937 (Table 7), Swift J1555.25402 (Table 8), 1RXS J170849.0400910 (Table 9), SGR 180620 (Table 10), XTE J1810197 (Table 11), Swift J1818.01607 (Table 12), Swift J1822.31606 (Table 13), SGR 18300645 (Table 14), 1E 1841045 (Table 15), SGR 1935+2154 (Table 16), 1E 2259+586 (Table 17), and the magnetar-like pulsar PSR J18460258 (Table 18). For magnetars with more than 30 detected bursts, a complete burst catalog is provided as an online table.
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 5505010502 | 2022-07-24T23:59:07.273 | 0.005 | |||||
| 2 | 5505011602 | 2023-06-24T12:28:23.950 | 795.595 | |||||
| 3 | 5505011703 | 2023-07-30T22:40:23.300 | 250.465 | 1.2 | 0.8 | |||
| 4 | 5505011704 | 2023-07-31T20:21:10.490 | 1055.592 | |||||
| ∗ Burst for which could not be constrained during spectral fitting is marked with “” | ||||||||
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 0020040102 | 2017-07-15T14:15:37.947 | 27.684 | |||||
| 2 | 2598061601 | 2019-10-20T00:30:01.984 | 10.171 | 1.2 | 0.8 | |||
| 3 | 4549061701 | 2021-10-22T01:41:59.761 | 19.641 | 1.2 | 0.8 | |||
| 4 | 6020040102 | 2023-09-26T11:56:38.889 | 287.419 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 1020240141 | 2018-11-22T11:58:14.425 | 0.001 | 1.2 | 0.8 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 4202190101 | 2021-06-03T13:58:12.407 | 15.382 | 1.2 | 0.8 | |||
| 2 | 4202190101 | 2021-06-03T13:58:30.247 | 68.397 | 1.2 | 0.8 | |||
| 3 | 4202190101 | 2021-06-03T14:01:43.339 | 18.493 | |||||
| 4 | 4202190101 | 2021-06-03T14:08:53.301 | 13.170 | 1.2 | 0.8 | |||
| 5 | 4202190101 | 2021-06-03T18:56:09.216 | 3.999 | 1.2 | 0.8 | |||
| 6 | 4560010101 | 2021-06-04T14:44:33.324 | 14.713 | 1.2 | 0.8 | |||
| 7 | 4560010101 | 2021-06-04T14:50:53.414 | 15.933 | 1.2 | 0.8 | |||
| 8 | 4560010102 | 2021-06-05T14:02:44.083 | 16.026 | 1.2 | 0.8 | |||
| 9 | 4560010102 | 2021-06-05T15:49:04.054 | 15.114 | 1.2 | 0.8 | |||
| 10 | 4560010102 | 2021-06-05T15:50:02.211 | 15.034 | 1.2 | 0.8 | |||
| 11 | 4560010102 | 2021-06-05T17:18:34.725 | 6.273 | 1.2 | 0.8 | |||
| 12 | 4560010102 | 2021-06-05T17:19:28.827 | 1.912 | 1.2 | 0.8 | |||
| 13 | 4560010102 | 2021-06-05T22:02:08.210 | 96.565 | 1.2 | 0.8 | |||
| 14 | 4560010102 | 2021-06-05T23:38:47.741 | 11.873 | 1.2 | 0.8 | |||
| 15 | 4560010105 | 2021-06-08T00:37:11.129 | 21.066 | 1.2 | 0.8 | |||
| 16 | 4560010105 | 2021-06-08T11:26:42.315 | 12.274 | 1.2 | 0.8 | |||
| 17 | 4560010105 | 2021-06-08T11:30:17.135 | 3.891 | 1.2 | 0.8 | |||
| 18 | 4560010201 | 2021-06-09T05:58:17.772 | 20.052 | 1.2 | 0.8 | |||
| 19 | 4560010201 | 2021-06-09T07:31:09.537 | 10.284 | 1.2 | 0.8 | |||
| 20 | 4560010201 | 2021-06-09T10:41:20.317 | 8.136 | 1.2 | 0.8 | |||
| 21 | 4560010201 | 2021-06-09T13:48:15.408 | 21.464 | 1.2 | 0.8 | |||
| 22 | 4560010201 | 2021-06-09T13:49:47.011 | 20.901 | 1.2 | 0.8 | |||
| 23 | 4560010202 | 2021-06-10T02:07:01.017 | 14.500 | 1.2 | 0.8 | |||
| 24 | 4560010202 | 2021-06-10T05:12:16.355 | 25.015 | 1.2 | 0.8 | |||
| 25 | 4560010301 | 2021-06-11T10:46:48.341 | 7.549 | 1.2 | 0.8 | |||
| 26 | 4560010602 | 2021-06-17T02:58:45.587 | 4.830 | |||||
| 27 | 4560010602 | 2021-06-17T03:00:05.180 | 44.434 | |||||
| 28 | 4560010602 | 2021-06-17T03:06:05.556 | 30.304 | 1.2 | 0.8 | |||
| 29 | 4560010701 | 2021-06-19T04:34:47.771 | 24.491 | 1.2 | 0.8 | |||
| 30 | 4560010701 | 2021-06-19T04:39:46.659 | 7.865 | 1.2 | 0.8 | |||
| 74 | 4560013017 | 2021-10-23T06:18:22.667 | 22.435 | 1.2 | 0.8 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 3622050401 | 2020-04-03T18:13:10.886 | 111.437 | |||||
| ∗ Burst for which could not be constrained during spectral fitting is marked with “” | ||||||||
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 5020410102 | 2023-02-25T00:23:44.944 | 136.532 | |||||
| 2 | 5020410102 | 2023-02-25T00:33:59.721 | 171.007 | |||||
| 3 | 5020410102 | 2023-02-25T01:32:03.158 | 163.288 | |||||
| 4 | 5020410103 | 2023-02-28T21:10:53.593 | 355.533 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 2020420105 | 2019-03-31T22:57:55.917 | 195.301 | |||||
| ∗ Burst for which could not be constrained during spectral fitting is marked with “” | ||||||||
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 3201060101 | 2020-03-13T04:47:22.249 | 17.981 | 1.2 | 0.8 | |||
| 2 | 3201060101 | 2020-03-13T06:06:49.313 | 94.302 | 1.2 | 0.8 | |||
| 3 | 3556010301 | 2020-03-19T18:42:30.401 | 32.664 | |||||
| 4 | 3556010401 | 2020-03-20T11:40:17.451 | 17.572 | 1.2 | 0.8 | |||
| 5 | 3556010401 | 2020-03-20T19:25:38.571 | 13.051 | 1.2 | 0.8 | |||
| 6 | 3556010501 | 2020-03-21T01:43:57.860 | 16.816 | 1.2 | 0.8 | |||
| 7 | 3556010501 | 2020-03-21T09:02:31.782 | 11.639 | 1.2 | 0.8 | |||
| 8 | 3556011501 | 2020-03-30T08:40:42.423 | 4.766 | 1.2 | 0.8 | |||
| 9 | 3556011502 | 2020-03-31T12:37:33.925 | 16.717 | 1.2 | 0.8 | |||
| 10 | 3556011502 | 2020-03-31T12:45:41.940 | 7.767 | 1.2 | 0.8 | |||
| 11 | 3556011601 | 2020-04-03T13:20:28.393 | 4.307 | 1.2 | 0.8 | |||
| 12 | 3556011801 | 2020-04-08T05:25:28.223 | 1.836 | 1.2 | 0.8 | |||
| 13 | 3556012001 | 2020-04-14T21:06:03.679 | 49.414 | 1.2 | 0.8 | |||
| 14 | 3556012301 | 2020-04-23T07:51:42.142 | 17.795 | 1.2 | 0.8 | |||
| 15 | 3556012401 | 2020-04-27T08:01:52.434 | 22.607 | |||||
| 16 | 3556012501 | 2020-04-29T22:01:37.866 | 64.451 | |||||
| 17 | 3556012602 | 2020-05-03T06:16:44.138 | 16.749 | 1.2 | 0.8 | |||
| 18 | 3556012602 | 2020-05-03T07:51:32.437 | 46.881 | 1.2 | 0.8 | |||
| 19 | 3556012602 | 2020-05-03T18:35:55.093 | 5.422 | 1.2 | 0.8 | |||
| 20 | 3556012701 | 2020-05-05T03:17:15.436 | 54.100 | 1.2 | 0.8 | |||
| 21 | 3556012901 | 2020-05-20T23:01:47.068 | 49.164 | 1.2 | 0.8 | |||
| 22 | 3556013401 | 2020-06-05T00:00:36.972 | 3.188 | 1.2 | 0.8 | |||
| 23 | 3556013701 | 2020-06-14T03:57:27.920 | 40.289 | |||||
| 24 | 3556013901 | 2020-06-20T10:39:12.226 | 14.965 | 1.2 | 0.8 | |||
| 25 | 3556014302 | 2020-07-02T02:31:37.178 | 3.395 | 1.2 | 0.8 | |||
| 26 | 3556014302 | 2020-07-02T05:35:54.337 | 28.858 | 1.2 | 0.8 | |||
| 27 | 3556015201 | 2020-08-25T09:48:03.712 | 53.695 | 1.2 | 0.8 | |||
| ∗ Burst for which could not be constrained during spectral fitting is marked with “” | ||||||||
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 1020460110 | 2018-04-11T19:53:02.682 | 206.345 | |||||
| ∗ Burst for which could not be constrained during spectral fitting is marked with “” | ||||||||
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 3201810102 | 2020-10-11T13:13:22.627 | 14.583 | 1.2 | 0.8 | |||
| 2 | 3201810103 | 2020-10-12T03:08:29.432 | 23.174 | 1.2 | 0.8 | |||
| 3 | 3201810103 | 2020-10-12T11:06:46.629 | 15.462 | 1.2 | 0.8 | |||
| 4 | 3201810104 | 2020-10-13T02:24:26.198 | 11.321 | 1.2 | 0.8 | |||
| 5 | 3201810104 | 2020-10-13T02:25:43.723 | 27.985 | 1.2 | 0.8 | |||
| 6 | 3201810104 | 2020-10-13T04:12:47.104 | 43.532 | |||||
| 7 | 3201810104 | 2020-10-13T08:51:13.440 | 31.461 | 1.2 | 0.8 | |||
| 8 | 3201810105 | 2020-10-14T05:53:25.046 | 17.298 | 1.2 | 0.8 | |||
| 9 | 3201810106 | 2020-10-17T12:51:50.936 | 55.452 | |||||
| 10 | 3201810106 | 2020-10-17T13:02:25.535 | 18.565 | 1.2 | 0.8 | |||
| 11 | 3201810106 | 2020-10-17T17:32:44.182 | 44.246 | 1.2 | 0.8 | |||
| 12 | 3201810106 | 2020-10-17T19:03:51.189 | 11.280 | 1.2 | 0.8 | |||
| 13 | 3201810106 | 2020-10-17T19:05:45.315 | 6.752 | 1.2 | 0.8 | |||
| 14 | 3201810106 | 2020-10-17T20:39:40.869 | 6.332 | 1.2 | 0.8 | |||
| 15 | 3201810106 | 2020-10-17T20:40:22.716 | 21.392 | 1.2 | 0.8 | |||
| 16 | 3201810106 | 2020-10-17T20:44:10.104 | 33.586 | |||||
| 17 | 3201810106 | 2020-10-17T20:44:43.164 | 28.339 | 1.2 | 0.8 | |||
| 18 | 3201810106 | 2020-10-17T20:47:13.721 | 59.941 | 1.2 | 0.8 | |||
| 19 | 3201810106 | 2020-10-17T20:49:17.019 | 22.958 | 1.2 | 0.8 | |||
| 20 | 3201810107 | 2020-10-18T05:52:55.466 | 11.258 | 1.2 | 0.8 | |||
| 21 | 3201810107 | 2020-10-18T06:03:31.423 | 41.537 | 1.2 | 0.8 | |||
| 22 | 3201810107 | 2020-10-18T07:38:29.114 | 17.312 | 1.2 | 0.8 | |||
| 23 | 3201810107 | 2020-10-18T07:40:44.267 | 24.162 | 1.2 | 0.8 | |||
| 24 | 3201810108 | 2020-10-19T03:44:59.635 | 17.771 | 1.2 | 0.8 | |||
| 25 | 3201810108 | 2020-10-19T06:47:05.411 | 4.292 | 1.2 | 0.8 | |||
| 26 | 3201810108 | 2020-10-19T09:46:53.410 | 4.465 | 1.2 | 0.8 | |||
| 27 | 3201810108 | 2020-10-19T09:56:39.345 | 27.231 | |||||
| 28 | 3201810108 | 2020-10-19T15:58:57.143 | 27.155 | 1.2 | 0.8 | |||
| 29 | 3201810109 | 2020-10-20T23:05:23.906 | 55.510 | |||||
| 30 | 3201810109 | 2020-10-20T23:08:31.159 | 49.872 | 1.2 | 0.8 | |||
| 128 | 4201810109 | 2021-03-22T05:28:50.886 | 28.740 | 1.2 | 0.8 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 7020500115 | 2024-08-21T05:47:16.995 | 275.708 | 1.2 | 0.8 | |||
| 2 | 7020500115 | 2024-08-21T05:53:42.996 | 125.542 | 1.2 | 0.8 | |||
| 3 | 7020500115 | 2024-08-21T05:55:46.567 | 1563.466 | |||||
| 4 | 7020500115 | 2024-08-21T07:24:50.450 | 778.175 | |||||
| 5 | 7020500115 | 2024-08-21T07:34:23.033 | 933.788 | |||||
| 6 | 7020500115 | 2024-08-21T07:38:19.847 | 73.198 | 1.2 | 0.8 | |||
| 7 | 7020500115 | 2024-08-21T07:38:52.599 | 134.445 | |||||
| 8 | 7020500115 | 2024-08-21T10:41:11.236 | 392.833 | |||||
| 9 | 7020500115 | 2024-08-21T10:44:00.142 | 369.713 | |||||
| 10 | 7020500115 | 2024-08-21T10:44:00.976 | 195.195 | 1.2 | 0.8 | |||
| 11 | 7020500116 | 2024-08-22T08:18:31.181 | 200.954 | |||||
| 12 | 7020500119 | 2024-08-25T00:05:16.059 | 42.214 | |||||
| 13 | 7020500133 | 2024-09-25T16:42:06.956 | 80.832 | 1.2 | 0.8 | |||
| 14 | 7020500136 | 2024-09-28T19:00:50.111 | 20.617 | 1.2 | 0.8 | |||
| 15 | 7582010112 | 2024-11-09T19:06:31.623 | 82.960 | 1.2 | 0.8 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 3020560101 | 2020-04-28T00:42:04.992 | 200.675 | |||||
| 2 | 3020560101 | 2020-04-28T00:42:07.375 | 308.612 | |||||
| 3 | 3020560101 | 2020-04-28T00:42:09.715 | 89.207 | 1.2 | 0.8 | |||
| 4 | 3020560101 | 2020-04-28T00:42:15.538 | 422.493 | |||||
| 5 | 3020560101 | 2020-04-28T00:42:32.978 | 641.929 | |||||
| 6 | 3020560101 | 2020-04-28T00:42:37.112 | 58.988 | 1.2 | 0.8 | |||
| 7 | 3020560101 | 2020-04-28T00:42:40.065 | 1452.321 | |||||
| 8 | 3020560101 | 2020-04-28T00:42:43.945 | 97.369 | |||||
| 9 | 3020560101 | 2020-04-28T00:42:44.229 | 48.001 | 1.2 | 0.8 | |||
| 10 | 3020560101 | 2020-04-28T00:42:45.235 | 569.746 | |||||
| 11 | 3020560101 | 2020-04-28T00:42:52.979 | 166.817 | 1.2 | 0.8 | |||
| 12 | 3020560101 | 2020-04-28T00:42:57.998 | 472.229 | |||||
| 13 | 3020560101 | 2020-04-28T00:43:05.677 | 25.986 | 1.2 | 0.8 | |||
| 14 | 3020560101 | 2020-04-28T00:43:10.501 | 509.074 | |||||
| 15 | 3020560101 | 2020-04-28T00:43:27.446 | 222.850 | |||||
| 16 | 3020560101 | 2020-04-28T00:43:29.533 | 234.156 | 1.2 | 0.8 | |||
| 17 | 3020560101 | 2020-04-28T00:43:31.919 | 327.903 | |||||
| 18 | 3020560101 | 2020-04-28T00:43:35.627 | 998.136 | |||||
| 19 | 3020560101 | 2020-04-28T00:43:38.163 | 355.770 | |||||
| 20 | 3020560101 | 2020-04-28T00:43:45.520 | 955.644 | |||||
| 21 | 3020560101 | 2020-04-28T00:43:52.985 | 1904.002 | |||||
| 22 | 3020560101 | 2020-04-28T00:44:00.547 | 445.025 | |||||
| 23 | 3020560101 | 2020-04-28T00:44:06.233 | 235.896 | |||||
| 24 | 3020560101 | 2020-04-28T00:44:07.030 | 183.750 | |||||
| 25 | 3020560101 | 2020-04-28T00:44:07.493 | 334.338 | |||||
| 26 | 3020560101 | 2020-04-28T00:44:08.441 | 735.361 | |||||
| 27 | 3020560101 | 2020-04-28T00:44:10.635 | 120.500 | |||||
| 28 | 3020560101 | 2020-04-28T00:44:17.045 | 901.641 | |||||
| 29 | 3020560101 | 2020-04-28T00:44:19.704 | 425.656 | |||||
| 30 | 3020560101 | 2020-04-28T00:44:20.470 | 56.817 | |||||
| 867 | 5576010116 | 2022-11-15T11:48:59.834 | 222.401 | 1.2 | 0.8 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 7020600108 | 2024-12-27T21:14:02.224 | 236.260 | 1.2 | 0.8 |
| Burst | ObsID | TDB start | HR | |||||
| (ms) | (keV) | (erg cm-2 s-1) | (erg cm-2 s-1) | |||||
| 1 | 3033290103 | 2020-08-05T17:08:55.807 | 20.480 | 1.2 | 0.8 | |||
| 2 | 3033290104 | 2020-08-06T22:25:26.265 | 32.558 | |||||
| 3 | 3033290108 | 2020-08-10T08:51:46.075 | 10.516 | 1.2 | 0.8 | |||
| 4 | 3598010801 | 2020-07-25T19:38:03.338 | 11.285 | 1.2 | 0.8 |